OpenStudy (anonymous):

4 fair coins are tossed. What is the probability that all the outcomes are heads?

OpenStudy (anonymous):

1/16

OpenStudy (shadowlegendx):

The probability that one coin will be heads= 1/4 The probability that two coins will be heads= 1/8 The probability that three coins will be heads= 1/12 The probability that four coins will be heads= 1/16 That should help you understand.

OpenStudy (anonymous):

thank you so much!!!!!

OpenStudy (shadowlegendx):

No Problem :)

hero (hero):

Actually The probability that one coin will be heads= 1/2 The probability that two coins will be heads= 1/4 The probability that three coins will be heads= 1/8 The probability that four coins will be heads= 1/16

OpenStudy (shadowlegendx):

I knew I did something funky in there, thanks for pointing it out @Hero

hero (hero):

You're welcome